Ticket Number: HEA-NOC/20040714-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: DIT
Ticket Opened: 20040714 07:15 UTC+1 Problem Start: 20040714 05:56 UTC+1
Ticket Update: 20040714 12:33 UTC+1 Problem End: 20040714 11:00 UTC+1
DIT access circuit, AVC502353, goes down.
Loss of connectivity to DIT. Mail backup via NUIM.
20040714 07:06 MN Check routes; only MAN pilot subnet visible. Report
fault to Eircom ATM faults; logged as #912.
20040714 07:25 MN Eircom phone to say they cannot see the ACE CPE at the
DIT end of the circuit. Will make contact with DIT personnel.
20040714 08:23 MN Dit personnel investigating.
20040714 08:25 MN DIT report loss of power in comms room.
20040714 10:03 RD DIT primary and backup endpoints are now responsive to
ping's. However no general traffic is flowing in or out of DIT yet - DIT
personnel are investigating further. DIT have been informed of necessary
backup MX change.
20040714 12:25 RD Traffic was not flowing due to a firewall problem, which
has been corrected. Normal traffic is flowing again from 11:00. The IPv6
tunnel is still non-responsive.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20040714-1